[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat May 06, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat May 6 18:43:47 UTC 2017


HSU Shift Log for Sat May 06 2017 14:43:38 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: It is super cold in operations. I can barely feel  
my hands.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170506/98089b26/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list